a college football game generated $85,732.50 on the sale of 14,910 tickets if student tickets sold for $4.50 and non-student tickets were sold for $9.75 how many of each type of ticket were sold
Answer by ReadingBoosters(3246)

Setting the variables
x = student tickets
y = non-student tickets
...
Given
x+y=14910
4.5x + 9.75y = $85,732.50
...
Substitute x into the second equation
4.5(14910-y) + 9.75y = 85732.50
67095 - 4.5y + 9.75y = 85732.50
Combine like terms
5.25y = 85732.50 - 67095
5.25y = 18637.50
Divide both sides by 5.25
y = 3,550 non-student tickets sold
x = 14910 - 3550 = 11360 student tickets sold
...
Check
4.5(11360) + 9.75(3550) = 51120 + 34612.5 = 85732.50
.........................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.
www.MyHomeworkAnswers.com